When a 60-kg boy sits at rest on a trampoline, it says 0.10 m at the center.

a. What is the effective spring constant for the trampoline?

b. The trampoline is pulled downward an extra 0.050 m by a strap sewed under the center of the trampoline. When the strap is released, what are the energy and frequency of the trampoline?
Answer #1

a)

Here force due to gravity is balanced by spring restoring force

mg =Kx

K=mg/x =60*9.81/0.1

K=5886 N/m

b)

Energy of the trampoline

E=(1/2)Kx2 = (1/2)*5886*0.052

E=7.358 J

frequency

f=(1/2pi)sqrt[K/m]

f=(1/2pi)sqrt[5886/60]

f=1.576 Hz